# Distill-SynthKG: Distilling Knowledge Graph Synthesis Workflow for Improved Coverage and Efficiency

### v1

@@ Appendix A LLM Prompts for SynthKG
Extract all named entities from the document. Also generate the type for each entity.
Instructions
- Generate only the most informative name for each named entity. Example: if John P., Parker, John Parker are coreferential, only generate John Parker.
@@
Extract all facts from the document. For each fact, also generate all semantic triplets.
@@
"fact": "Mounjaro generated $2.2 billion in sales in the fourth quarter of 2023, an eight-fold increase from the $279 million during the same period
in 2022.",
"triplets": [["Mounjaro", "2023’s fourth quarter sales", "$2.2 billion sales"], ["Mounjaro", "2022’s fourth quarter sales", "$279 million"]]
@@
"fact": "Eli Lilly’s fourth-quarter sales were around $9.4 billion, a 28% increase over the $7.3 billion during the same period in 2022.",
"triplets": [["Eli Lilly", "2023 fourth-quarter sales", "$9.4 billion,"], ["Eli Lilly", "2022 fourth-quarter sales", "$7.3 billion,"]]
Appendix B KG Coverage Evaluation
B.1 Prompts and examples
@@
You are given a question answer pair, please generate a relation triplet to represent the relationship.
Generate output in the format described below. “‘ head || relation || tail “‘ Note: - Must include relation, head entity and tail entity. Ensure that head entity is a subject of relation, and tail entity is a direct object of relation. - You must use the given answer as the head or tail entity. - Specific entity is more preferable than generic entity. - Do NOT generate pronouns or references in head and tail entities. —- Example 1:
Question: To whom was Messi’s goal in the first leg of the Copa del Rey compared? Answer: Diego Maradona
Output: Messi’s goal || was compared to || Diego Maradona
Example 2:
Question: The father of Chiang Hsiao-wen is whom? Answer: Chiang Ching-kuo
Output: Chiang Ching-kuo || the father of || Chiang Hsiao-wen
Question: {question}
Answer: {answer}
Output:
You are given a multihop question, some facts that are used to reach the correct answer, and the correct answer. Your goal is to decompose the question into sub-question, and the corresponding answer to each sub question.
Example 1
Question: What relationship does Fred Gehrke have to the 23rd overall pick in the 2010 Major League Baseball Draft?
Facts: He is the great-grandfather of Miami Marlin Christian Yelich Yelich was drafted out of high school by the Marlins in the 1st round (23rd overall) of the 2010 Major League Baseball Draft.
Answer: great-grandfather
Decompose question answer pairs:
Who was the 23rd overall pick in the 2010 Major League Baseball Draft? Christian Yelich
What relationship does Fred Gehrke have to Christian Yelich? Great-grandfather
Question: {question}
Facts: {facts}
Answer: {answer}
Decompose question answer pairs:
B.2 Human evaluation of extracted triplets
@@
Appendix C Experimental Settings
C.1 QLoRa fine-tuning setup
In our experiments detailed in Section 7.5, we employ the QLoRA fine-tuning. The training configuration used is as follows: we train models for 3 epochs, with an alpha value of 256 and a rank of 128. The learning rate, warmup steps and batch size are set to 0.00003, 50 and 8 respectively.
Appendix D Task-Specific Evaluation Settings
D.1 KG Coverage Task
@@
D.2 Retrieval Task
We use ‘text-embedding-3-small’ for both dense retrieval and embedding propositions in KG-based retrievers. For both the graph and graph + LLM retrievers, we first construct the sub-graph by selecting the 200 propositions most similar to the question based on embedding similarity. Within the sub-graph, we traverse the KG, starting from the question entity, and select propositions within a 5-hop neighborhood. For re-ranking the propositions in the LLM-based retriever and also for re-ranking chunks in Dense + LLM retriever, we use the GPT-4o model. The Dense+LLM retriever uses LlamaIndex’s implementation of the LLMRerank post-processor.
D.3 Multi-hop Question Answering Task
D.3.1 LlamaIndex Configuration
Table 6 presents the complete configuration of our LlamaIndex query engine setup.
D.3.2 Chain-of-Triplet
We design a triplet retrieval method that first breaks down the question into sub-queries in a triplet format. These triplet queries are then used to retrieve the most semantically matching triplet facts from the extracted knowledge graph. Specifically, it includes three steps to generate the final answer.
@@ D.3.2 Chain-of-Triplet
given a question, we convert it into a series of triplet queries. Specifically, since our downstream task involves multi-hop question answering, instead of generating a single triplet, we prompt the model to generate a chain of triplets. The generated triplets may contain placeholders that represent unknown entities. The prompt is shown in Figure 12.
There is a knowledge graph (entities and relations). Now, you are given a question, your task is to decompose this question into a chain of triplets used for searching fact from the graph.
The triplet should be in this format: head || relation || tail
Note: - Ensure that head entity is a subject of relation, and tail entity is a direct object of relation. - Do NOT generate pronouns or references in head and tail entities. - Do NOT generate entities that are not appeared in the question. - If an triplet includes an intermediate answer or the final answer, you can use # followed by an digit for reference. - The triplets order should be the same order for retrieving the facts from a knowledge graph.
Example 1:
Question: Who is older, Hampton Del Ruth or Ted Kotcheff?
Decompose Triplets:
Hampton Del Ruth || was born on || #1 Ted Kotcheff || was born on || #2
Example 2:
Question: In what town is Suffolk county hamlet that was served by the Suffolk Traction Company?
Decompose Triplets:
Suffolk Traction Company || served || #1 #1 || is located in || #2
Now please generate the decompose Triplets for the question: {question}
Decompose Triplets:
@@
You are given a natural language question, triplets chain for this question, a set of retrieved tripletes, and a set of facts, please answer the question.
Question: {question}
Question Triplets Chain: {question triplets}
Retrieved Triplets: {retrieved triplets}
Retrieved Facts: {retrieved facts}
Short Answer no more than 3 words:
@@
: We use the same graph + LLM retriever hyper-parameters as in section D.2.
@@
Appendix E License information

### v2

@@ Appendix C LLM Prompts for SynthKG
"fact": "Mounjaro generated $2.2 billion in sales in the fourth quarter of 2023,
an eight-fold increase from the $279 million during the same period in 2022.",
"triplets": [["Mounjaro", "2023’s fourth quarter sales", "$2.2 billion sales"],
["Mounjaro", "2022’s fourth quarter sales", "$279 million"]]
@@
"fact": "Eli Lilly’s fourth-quarter sales were around $9.4 billion,
a 28% increase over the $7.3 billion during the same period in 2022.",
"triplets": [["Eli Lilly", "2023 fourth-quarter sales", "$9.4 billion,"],
["Eli Lilly", "2022 fourth-quarter sales", "$7.3 billion,"]]
Appendix D KG Coverage Evaluation
D.1 Prompts and examples
@@
D.2 Human evaluation of extracted triplets
@@
Appendix E Experimental Settings
E.1 QLoRa fine-tuning setup
In our experiments detailed in Section˜B.1, we employ the QLoRA fine-tuning. The training configuration used is as follows: we train models for 3 epochs, with an alpha value of 256 and a rank of 128. The learning rate, warmup steps and batch size are set to 0.00003, 50 and 8 respectively.
Appendix F Task-Specific Evaluation Settings
F.1 KG Coverage Task
@@
F.2 Retrieval Task
We use ‘text-embedding-3-small’ for both dense retrieval and embedding propositions in KG-based retrievers. For both the graph and graph + LLM retrievers, we first construct the sub-graph by selecting the 200 propositions (M = 200) most similar to the question based on embedding similarity. Within the sub-graph, we traverse the KG, starting from the question entity, and select propositions within a 5-hop neighborhood (N = 5). For re-ranking the propositions in the LLM-based retriever and also for re-ranking chunks in Dense+LLM retriever, we use the GPT-4o model. The Dense+LLM retriever uses LlamaIndex’s implementation of the LLMRerank post-processor.
We evaluate retrieval performance at the passage level, following the setup used in HippoRAG . For each query, we evaluate whether the retrieved passages contain all ground-truth information required to answer the question. Retrieval metrics include Hits@2 and Mean Reciprocal Rank (MRR), computed based on the rank positions of the passages associated with ground-truth triplets. Specifically, a passage is considered relevant if it contains all the facts (triplets or propositions) needed for correct multi-hop reasoning. The evaluation code is directly adopted from HippoRAG.
To ensure comparability, we use the same benchmark datasets and experimental setup as HippoRAG, including 1,000 questions and a fixed set of candidate passages. The number of ground-truth passages per question is consistent with the original annotations. Our experiments are conducted on three multi-hop QA datasets: MuSiQue (11,656 passages), 2WikiMultiHopQA (6,119 passages), and HotpotQA (9,221 passages).
F.3 Multi-hop Question Answering Task
F.3.1 LlamaIndex Configuration
Table˜7 presents the complete configuration of our LlamaIndex query engine setup.
F.3.2 Chain-of-Triplet
We design a triplet retrieval method that first breaks down the question into sub-queries in a triplet format. These triplet queries are then used to retrieve the most semantically matching triplet facts from the extracted KG. Specifically, it includes three steps to generate the final answer.
@@ F.3.2 Chain-of-Triplet
given a question, we convert it into a series of triplet queries. Specifically, since our downstream task involves multi-hop QA, instead of generating a single triplet, we prompt the model to generate a chain of triplets. The generated triplets may contain placeholders that represent unknown entities. The prompt is shown in Figure 11.
@@
: We use the same graph + LLM retriever hyper-parameters as in section˜F.2.
Appendix G Data Release and Training/ Inference Cost Considerations
We will make our annotated 100K data samples publicly available to support future research. With the rapid advancements in LLMs, researchers may choose to resynthesize data to better align with their specific applications. In such cases, we recommend using our cost-efficient approach, detailed in Section˜B.1, which provides a practical balance between performance and computational cost.
Below, we present the detailed training and inference costs, highlighting the efficiency of our SynthKG and Distill-SynthKG methods.
Cost of Data Synthesis:
With the Llama-3.1-70b-Instruct model, running the entire SynthKG pipeline on a single document requires processing an average of 11,849 input tokens. This results in a total of 2,675 average output tokens, distributed across intermediate steps such as decontextualization and the final entities, relations, and proposition generation. At a cost of $0.90 per million tokens333https://www.together.ai/pricing, the total annotation cost per document is $0.0131. This translates to $13.08 for synthesizing training data for the model and $392.28 for the D-SynthKG-8b model.
Cost of Model Training:
After consolidating the data synthesized by SynthKG, each document contains an average of 1,723 input tokens (including prompts) and 1,248 output tokens, totaling 2,971 tokens per document. For a dataset of 30,000 documents, the total training token count is 89.13 million tokens. Fine-tuning Llama-3.1-8b-Instruct for one epoch on this dataset to obtain D-SynthKG-8b would cost $36.65. Additionally, fine-tuning Mistral-7b-Instruct-v0.2 for 3 epochs to obtain the model would cost $3.67.
Combining data synthesis and fine-tuning costs, training D-SynthKG-8b would cost $428.93, while training would cost $16.75.
Inference Cost:
As mentioned in the cost of data synthesis, processing a single document using the SynthKG pipeline requires an average of 11,849 input tokens and 2,675 output tokens, totaling 14,524 tokens per document. At a cost of $0.90 per million tokens, this amounts to $0.031 per document. In contrast, with D-SynthKG-8b and , each document requires 1,723 input tokens and 1,248 output tokens, totaling 2,971 tokens, with a cost of $0.00267 per document. This is only 8.6% of the cost of SynthKG, demonstrating the significant efficiency gains from fine-tuning a distilled model.
@@ Appendix G Data Release and Training/ Inference Cost Considerations
Appendix H License information
@@
Appendix I Discussion on the Segmentation Strategy and Document Structure Preservation
While our segmentation and de-semanticization approach may appear simple, our design choice is guided by both empirical findings and practical considerations. First, our analysis (see Figure 4) shows that model performance consistently degrades as document length increases. Preserving structural associations in segmentation might result in longer input spans, which, based on our experiments, would still harm performance. Second, while we agree that capturing structural dependencies across multiple pages is a meaningful goal, it remains an open research challenge—particularly in open-ontology settings. Even state-of-the-art models like GPT-4o lack robust, generalizable pipelines for reliably preserving cross-page structure in a way that could be distilled into a smaller model. Given these limitations, we prioritize practicality and scalability by adopting a fixed-size ( 1K token) chunking approach. This method aligns with the constraints of retrieval-augmented generation (RAG) and enables effective, document-level KG construction at scale. We believe our approach provides a strong balance between empirical performance and real-world applicability under current technological constraints.
Appendix J Discussion on Manual Hyperparameter Selection
We manually tune two key hyperparameters in our framework: the semantic match score threshold for triplet coverage evaluation and the ROUGE score threshold for decontextualization. For the semantic match score, we intentionally select a higher threshold to ensure accuracy when determining whether two triplets are semantically equivalent. It is important to note that this threshold is only used for evaluation purposes and does not affect model training, inference, or RAG evaluation, thus having no impact on the model’s learning or predictions. While the threshold is manually adjusted, downstream task performance reflects the reliability of our model.
For the ROUGE score threshold used in decontextualization, we conduct careful manual analysis to ensure its effectiveness. For both Llama-3.1-70b and GPT-4o, we examine a small subset of chunks with low ROUGE scores and find that most rewritings are accurate. We identify 593 edits in this subset, with only six containing incorrect modifications and four showing some loss of information. These results suggest that the decontextualization process generally yields high-quality, self-contained text.
The low ROUGE scores typically result from document footers or metadata, which are removed during the LLM’s decontextualization process, and not from factual errors or information loss. Our analysis shows that approximately 72% of the chunks achieve a ROUGE score of 90 or higher, reflecting strong alignment with the original content. Chunks with scores below 0.70 make up less than 3%, and these are filtered out during the process to avoid any omission or extreme paraphrasing.
Appendix K LLM Usage Statement
We used ChatGPT and Claude as writing assistants to improve the clarity and grammar of our manuscript. All scientific content, experimental design, and technical contributions are the original work of the authors.